### v1

# Constrained Bandwidth Observation Sharing for Multi-Robot Navigation in Dynamic Environments via Intelligent Knapsack

Abstract
Multi-robot navigation is increasingly crucial in various domains, including disaster response, autonomous vehicles, and warehouse and manufacturing automation. Robot teams often must operate in highly dynamic environments and under strict bandwidth constraints imposed by communication infrastructure, rendering effective observation sharing within the system a challenging problem. This paper presents a novel optimal communication scheme, Intelligent Knapsack (iKnap), for multi-robot navigation in dynamic environments under bandwidth constraints. We model multi-robot communication as belief propagation in a graph of inferential agents. We then formulate the combinatorial optimization for observation sharing as a 0/1 knapsack problem, where each potential pairwise communication between robots is assigned a decision-making utility to be weighed against its bandwidth cost, and the system has some cumulative bandwidth limit. Compared to state-of-the-art broadcast-based optimal communication schemes, iKnap yields significant improvements in navigation performance with respect to scenario complexity while maintaining a similar runtime. Furthermore, iKnap utilizes allocated bandwidth and observational resources more efficiently than existing approaches, especially in very low-resource and high-uncertainty settings. Based on these results, we claim that the proposed method enables more robust collaboration for multi-robot teams in real-world navigation problems.
I Introduction
The rapid proliferation of multi-robot systems across various domains has underscored the critical importance of efficient communication strategies to navigate the environment and coordinate actions [1]. Autonomous vehicles can leverage over-the-air communication to share information about the road environment, yielding safer transportation systems [2, 3]. Search-and-rescue robot teams deployed in hazardous areas must efficiently relay real-time observations to locate survivors and assess damage [4]. In industrial environments such as manufacturing and warehouses, robot groups coordinate complex tasks such as material handling and inventory management [5, 6].
Much of the current work in multi-robot coordination assumes perfect communication within the system. However, in all the aforementioned settings, bandwidth limitations imposed by centralized or decentralized communication infrastructure are inherent and inevitable [7]. Moreover, the increasingly complex and dynamic nature of the tasks and environments in which multi-robot systems are deployed will demand greater levels of coordination, and thus more communication [1]. This necessitates the treatment of bandwidth as a luxury rather than a guarantee to ensure the feasibility of real-world implementation of these systems. Thus, we are motivated to pursue utility-conscious communication schemes for robot teams.
Figure 1: As robots $\mathcal{A}_{1}$ and $\mathcal{A}_{2}$ navigate through a warehouse environment, $\mathcal{A}_{1}$ observes moving obstacle $\mathcal{S}_{1}$ and shares its observation with $\mathcal{A}_{2}$ over the air. Consequently, $\mathcal{A}_{2}$ is made aware of a previously invisible potential collision with $\mathcal{S}_{1}$ and can plan accordingly.
There has been significant prior work on decision-theoretic approaches for communication in multi-agent systems. Many methods evaluate potential communications based on their expected effect on team reward [8, 9]. However, computing this value directly for general multi-agent models like Dec-MDPs or Dec-POMDPs is prohibitively complex [11, 10]. The complexity arises from the need to reason about the effects of communication on future team behavior, which requires modeling the beliefs and potential actions of all agents. Most existing methods attempting to solve this problem focus primarily on deciding the frequency of communication, e.g. via expected value techniques, assuming messages contain an agent’s entire observation history [12, 13]. However, this is impractical under real-world bandwidth constraints, where robots must carefully select which observations are most valuable to share.
To address the problem of minimizing the actual content of messages subject to bandwidth constraints, some initial work [14, 15] used greedy optimization to select observations to include in messages. While efficient, the method suffers poor scalability to more complex problems due to quickly decaying solution quality. More recent approaches have also explored using deep reinforcement learning to learn communication policies [16, 17]. While capable of scalability to very large scenarios, such learning-based approaches are usually not generalizable to a broader class of multi-robot navigation problems without extensive training. Thus, prioritizing generalizability over radical scalability, we are motivated to pursue an optimization-based approach for determining communications.
Most recently, Marcotte et al. [18] proposed a method for optimizing multi-robot communication under bandwidth constraints using a bandit-based combinatorial optimization technique to select observations for inclusion in messages. However, the formulation relies on a broadcast-based communication scheme where messages are shared with all agents simultaneously. In practice, the utility of a particular piece of information for decision-making is often highly subjective to each individual agent. For example, the observation of some obstacle might only be useful to agents whose planned paths potentially run into that obstacle, and irrelevant to those far away. In that case, by sending messages only to those who need them, we can save communication bandwidth as well as computation power without sacrificing performance. Thus, we are motivated to formalize the above intuition into a pairwise communication scheme.
In this paper, we present Intelligent Knapsack (iKnap), a novel optimal communication scheme for multi-robot observation sharing during navigation in dynamic environments under bandwidth constraints. Our method builds on insights from inferential statistics and combinatorial optimization to develop a principled yet tractable solution suitable for real-world implementation. Our key contributions are:
  • Modeling pairwise communication in a multi-robot system as graph-based multi-agent inference.
  • Formulation of the optimal communication problem as a 0/1 knapsack problem, enabling direct and efficient combinatorial optimization of the tradeoff between information value and communication resource usage.
  • Comprehensive empirical analysis of the proposed method’s navigation performance in increasingly complex environments and adaptation to various resource levels and uncertainty conditions.
The rest of our paper is organized as follows. Section II formulates the limited bandwidth observation sharing problem. Section III presents our 0/1 knapsack approach to solving this problem. Section IV discusses the results of simulated experiments using our approach. Finally, Section V concludes our study and presents future directions.
II Problem Formulation
We are given a multi-robot system of $n$ agents navigating through an environment where each agent has some goal location it must travel toward. The environment contains $m$ subjects, i.e. dynamic obstacles, that the agents wish to observe in order to improve decision-making during their navigation task. The agents and subjects together operate in a state space $\mathcal{X}$, where $\boldsymbol{x}^{A}_{i}\in\mathcal{X}$ denotes the state of the $i$-th agent and $\boldsymbol{x}^{S}_{k}\in\mathcal{X}$ denotes the state of the $k$-th subject. Moreover, the $i$-th agent’s belief about the $k$-th subject’s state variable at time $t$ is modeled by a multivariate probability density function $\boldsymbol{\Phi}^{t}_{ik}:\mathcal{X}\mapsto\mathbb{R}$. We shall assume that all agents’ states are known by each other.
Each agent $i$ continually obtains observations of the states of the observable subjects in its environment. The observation $\boldsymbol{\mu}_{ik}^{t}$ of subject $k$’s state at time $t$ is corrupted with some inherent noise $\boldsymbol{\sigma}_{ik}^{t*}$, and agent $i$ self-assigns some perceived uncertainty $\boldsymbol{\sigma}_{ik}^{t}$ to the observation. This represents real-world perception systems well, as a robot is obviously incapable of assessing the exact measurement error of its own sensors. As an agent accumulates more observations about a subject, it updates its beliefs about that subject $\boldsymbol{\Phi}^{t}_{ik}$ accordingly.
Agents can share observations with each other over the air, i.e. at some time $t$ during navigation, agent $a$ can share the observation $(\boldsymbol{\mu}^{t}_{ah},\boldsymbol{\sigma}^{t}_{ah})$ regarding subject $h$’s state with agent $b$, thus updating $b$’s belief $\boldsymbol{\Phi}^{t}_{bh}$ accordingly. At any $t$, the multi-robot system has a total communication bandwidth limit of $B$. Let communication between $a$ and $b$ regarding subject $h$ at time $t$ occupy some bandwidth $\beta^{t}_{abh}$. Moreover, denote the set of all communications $(a,b,h)$ occurring at time $t$ as $\mathcal{C}^{t}_{*}:=\{(a_{1},b_{1},h_{1}),...,(a_{N*},b_{N*},h_{N*})\}$. Then we have $\sum_{(a,b,h)\in\mathcal{C}^{t}_{*}}\beta^{t}_{abh}\leq B$ for all $t$.
The agents aim to minimize some navigation cost function $J(\boldsymbol{x^{A}})$ as they travel toward their goals from time $t=0$ to $t=T$. The agents communicate observations with periodicity $\delta t$, and this communication strategy potentially improves navigation performance as agents are made more aware of their environment, e.g. collision avoidance. Thus, the agents must solve the following optimization problem for communication-enabled navigation:
$\min\hskip 5.0pt\sum_{t=0}^{T}J(\boldsymbol{x^{A}})\hskip 2.0pt$ (1)
[…]
$\forall t\in(0,\delta t,...,T-\delta t,T)$
In the following section, we will solve this problem by investigating the construction of $\mathcal{C}^{t}_{*}$ more closely.
III Intelligent Knapsack Approach
In this section, we present the the iKnap optimal communication scheme in detail. We first formulate a pairwise communication scheme for the multi-robot system. We then present the combinatorial optimization problem for selecting observations to share and detail solving this problem via 0/1 knapsack.
III-A Communication Infrastructure
For more efficient processing of optimal communications, we establish a centralized communication infrastructure $\mathcal{I}$ to allocate bandwidth in coordinating precise pairwise communications between agents, as opposed to a broadcast-based infrastructure. Each agent $i$ then periodically uploads its state $\boldsymbol{x}^{A}_{i}$, observations $(\boldsymbol{\mu}_{i},\boldsymbol{\sigma}_{i})$, and beliefs $\boldsymbol{\Phi}_{i}$ to $\mathcal{I}$, which can then relay this information to the other agents. We assume that this upload process occupies some fixed bandwidth. Then the limited remaining bandwidth $B$ on $\mathcal{I}$ is utilized for sharing observations between the agents. In particular, $B$ is an upper bound on the volume of total simultaneous communication occurring through $\mathcal{I}$ at any given timestep.
III-B Pairwise Communication Scheme
At each timestep $t$, say we have a set of $N$ potential communications $\mathcal{C}^{t}=\{C^{t}_{1},...,C^{t}_{N}\}$ to choose from, and some subset $\mathcal{C}^{t}_{*}$ is the set of communications to be actually executed by $\mathcal{I}$ at time $t$. Each potential communication $C_{i}:=(a_{i},b_{i},h_{i},\beta_{i},\theta_{i})$ is directed from agent $a_{i}$ to agent $b_{i}$, concerning subject $h_{i}$, and has an associated bandwidth cost $\beta_{i}$ and associated utility $\theta_{i}$. One may consider $\theta_{i}$ as a measure of how “useful” $a_{i}$’s observation is for $b_{i}$’s decision-making. While it is difficult to measure utility objectively, various principled heuristics have been introduced in the literature [19, 20]. For potential communication $i$ from agent $a$ to agent $b$ regarding subject $h$, the utility heuristic function is $\theta_{i}(\boldsymbol{x}^{A}_{b},\boldsymbol{x}^{S}_{h},\boldsymbol{\Phi}^{A}_{a},\boldsymbol{\Phi}^{A}_{b})$. Note that our method can be generalized to a decentralized formulation to satisfy individual robots’ bandwidth constraints, where the utility heuristic function is $\theta_{i}(\boldsymbol{x}^{A}_{b},\boldsymbol{x}^{S}_{h},\boldsymbol{\Phi}^{A}_{a})$, assuming that state communication between robots occupies a fixed bandwidth.
In practice, $\mathcal{C}$ might contain all pairs of agents $(a_{i},b_{i})$, or be restricted to only those certain pairs in the system which are able to communicate. Agent communicability can be represented graphically, hence transitivity may apply, i.e. if $a$ communicates with $b$ with bandwidth $\beta_{1}$ and $b$ communicates with $c$ with bandwidth $\beta_{2}$, then $a$ can communicate with $c$ indirectly with bandwidth $\beta_{1}+\beta_{2}$.
(a) Observation Uploading
(b) iKnap Pairwise Communication
(c) Broadcast-Based Scheme
Figure 2: Example scenario involving a robot team $A_{1},A_{2},A_{3},A_{4}$ and a dynamic obstacle subject $S_{1}$ only visible by $A_{1}$ due to a wall obstruction. There is an impending collision between $A_{3}$ and $S_{1}$. In (a), all agents routinely upload their states and beliefs to the infrastructure $\mathcal{I}$, and $A_{1}$ additionally uploads its observation of $S_{1}$ to $\mathcal{I}$. Following this process, the proposed pairwise communication scheme is presented in (b), where $A_{1}$’s observation is shared only with $A_{3}$ to enable safe re-planning around $S_{1}$ while maintaining efficient communication. Alternatively, (c) depicts the traditional broadcast-based communication scheme, where $A_{1}$’s observation is shared with the entire robot team, despite the information’s irrelevance to the decision-making of $A_{2}$ and $A_{3}$.
III-C Combinatorial Optimization Problem
It follows that at each timestep $t$, we must select some subset of $\mathcal{C}^{t}$ to construct $\mathcal{C}^{t}_{*}$ for communication across $\mathcal{I}$. We shall denote this selection as a boolean vector $\boldsymbol{\psi}$, where $\psi_{i}=1$ implies $C^{t}_{i}\in\mathcal{C}^{t}_{*}$ and $\psi_{i}=0$ implies $C_{i}\notin\mathcal{C}^{t}_{*}$. At any given timestep, we seek to find the optimal communication subset $\boldsymbol{\psi}^{*}$ that maximizes the cumulative utility of the communications across $\mathcal{I}$ while satisfying the bandwidth limit $B$. Thus, at each timestep we can run the following optimization on $\mathcal{I}$:
$\boldsymbol{\boldsymbol{\psi}}^{*}=\operatorname*{argmax}_{\boldsymbol{\psi}}\hskip 5.0pt\boldsymbol{\psi}^{T}\boldsymbol{\theta}$ (2)
s.t. $\boldsymbol{\psi}^{T}\boldsymbol{\beta}\leq B,$
$\boldsymbol{\psi}\in\{0,1\}^{N}$
where $\boldsymbol{\theta}:=\{\theta_{1},...,\theta_{N}\}$ and $\boldsymbol{\beta}:=\{\beta_{1},...,\beta_{N}\}$.
This problem takes the form of a 0/1 knapsack problem, which is solvable in pseudo-polynomial time via dynamic programming. In particular, the algorithm finds a subset of $N$ items, each with some weight and value, to place in a knapsack of weight capacity $B$ such that the cumulative value is maximized in $O(NB)$ time. In our case, $N$ is $O(n^{2})$, yielding a time complexity of $O(n^{2}B)$. For a depiction of the overall observation sharing process, see Algorithm 1. For an example of the observation uploading process and a comparison of pairwise and broadcast-based communication schemes, see 2.
IV Experimental Results
In this section, we evaluate the proposed method under various simulated conditions and compare its performance to the current state of the art.
IV-A Simulation Setup
We design a randomized Python simulation environment with an open field containing $n$ robots, $m$ dynamic obstacles as subjects of observation, and $w$ walls as static obstacles to maneuver around. Robots and subjects follow 2D double-integrator dynamics from starting locations to goal locations. Robots obtain noisy observations of visible subjects around them and update their beliefs accordingly. We also implement a centralized communication infrastructure that collects uploaded observations and beliefs from each agent, runs iKnap to compute optimal communications, and executes these communications by sending observations to the relevant agents at some fixed frequency.
Algorithm 1 Observation sharing
1:Input: Set of agents $\mathcal{A}=\{a_{1},a_{2},...,a_{n}\}$, set of subjects $\mathcal{S}=\{s_{1},s_{2},...,s_{m}\}$, bandwidth limit $B$, communication period $\delta t$
2:Output: Updated beliefs $\boldsymbol{\Phi}$ for each agent
3:for each time step $t=0,\delta t,2\delta t,...,T$ do
4:     for each agent $i\in\mathcal{A}$ do
5:         $\boldsymbol{\mu}_{ik}^{t}\leftarrow\textsc{ObtainObservations}(i,\mathcal{S},t)$
6:         $\boldsymbol{\Phi}_{ik}^{t}\leftarrow\textsc{UpdateBelief}(i,\boldsymbol{\mu}_{ik}^{t})$
7:     $\mathcal{C}^{t}\leftarrow\{\}$
8:     for each pair of agents $(a,b)\in\mathcal{A}\times\mathcal{A},a\neq b$ do
9:         for each subject $h\in\mathcal{S}$ do
10:              $\beta_{abh}\leftarrow\textsc{BandwidthCost}(a,b,h)$
11:              $\theta_{abh}\leftarrow\textsc{Utility}(a,b,h,\boldsymbol{x}^{A}_{b},\boldsymbol{x}^{S}_{h},\boldsymbol{\Phi}^{A}_{a},\boldsymbol{\Phi}^{A}_{b})$
12:              Push$(\mathcal{C}^{t},(a,b,h,\beta_{abh},\theta_{abh}))$
13:     $\mathcal{C}^{t}_{*}\leftarrow\textsc{SolveKnapsackProblem}(\mathcal{C}^{t},B)$
14:     for each selected communication $C_{abh}^{*}\in\mathcal{C}^{t}_{*}$ do
15:         ShareObservation($a,b,h$)
16:         $\boldsymbol{\Phi}_{bh}^{t}\leftarrow\textsc{UpdateBelief}(b,\boldsymbol{\mu}_{ah}^{t})$
IV-B Perception Design
At each timestep, each agent obtains a noisy observation of each visible subject’s state, i.e. position and velocity. We define a subject to be visible to an agent if no wall exists between the subject and the agent, which reflects real-world occlusion scenarios. Then, between agent $i$ and subject $k$, the observation $\boldsymbol{\mu}_{ik}$ is corrupted with Gaussian noise
$\boldsymbol{\sigma}_{ik}=\frac{\alpha}{||\boldsymbol{x}^{A}_{i}-\boldsymbol{x}^{S}_{k}||^{2}},$ (3)
for tuning parameter $\alpha$, i.e. observations are less accurate the further away an agent is from a subject. Each time agent $i$ receives a new observation of subject $k$’s state, it updates its beliefs $\boldsymbol{\Phi}_{ik}$ via Bayesian inference with Gaussian prior and likelihood and assuming constant velocity. An agent’s confidence in its observation is proportional to the perceived distance between itself and the subject.
IV-C Utility Heuristic Design
The utility heuristic must provide an estimate of how useful an observation from agent $a$ regarding subject $h$ would be to agent $b$. Intuitively, we must take into account
  • the improvement in $b$’s belief about $h$ would improve as a result of $a$’s observation, and
  • the relevance of knowledge about $h$’s state in $j$’s navigation task.
For the former criterion, we choose to measure the Kullback–Leibler divergence between the prior and posterior distributions of $\boldsymbol{\Phi}_{bh}$ for position and velocity, using $a$’s observation as the likelihood distribution; we denote this value $\kappa_{abh}$. For the latter criterion, we choose to forward-simulate the trajectories of $b$ and $h$, assuming a constant-velocity trajectory for $h$ based on $a$’s beliefs, and then measure the inverse-squared minimum distance between $b$ and $h$ across a horizon of $t_{H}$ seconds in the navigation task; we denote this value $\tau_{abh}$. Then, the utility for potential communication $C_{i}$ associated with $(a,b,h)$ is given by
$\theta_{i}=p_{1}\kappa_{abh}+p_{2}\tau_{abh}$ (4)
for tuning parameters $p_{1}$ and $p_{2}$ to normalize and determine the relative importance of each metric.
IV-D Planner and Controller Design
Each agent plans a straight-line path from its start location to its goal location that wraps around walls where necessary. By default, each agent has some target speed along its trajectory. However, each agent runs a trajectory predictor assuming constant velocity based on its beliefs about the state of nearby subjects, and if it perceives an impending collision with a subject, it sets its target speed to 00 at the location of the collision. The distance-to-collision computation returns a probability distribution propogated from the uncertainty in the subject’s state, and the agent subtracts one standard deviation from the mean distance in order to ensure safety. The agent uses PD control to compute the corresponding acceleration at each timestep. Notice that this implementation rewards an agent’s greater confidence in its observations of subjects, as this directly improves navigation outcomes.
For each subject, we generate a randomized cubic spline trajectories from its start location to its goal location. We set some random target speed for the subject along its trajectory, and the subject uses PD control to compute the corresponding acceleration at each timestep.
IV-E Baseline and Evaluation Metrics
We compare the proposed framework’s performance to that of OCBC, a state-of-the-art observation sharing method proposed in [18]. Most notably, OCBC operates under a broadcast-based communication scheme, as opposed to a pairwise scheme. Additionally, OCBC leverages a bandit-based combinatorial optimization approach for selecting observations to share, as opposed to a knapsack approach. Consequently, OCBC is unable to account for variations in bandwidth required for each potential communication.
We measure algorithm performance via runtime where necessary, and we measure navigation performance via makespan, which is justified from the agent controller design presented in Section IV-D. All experiments were run on a desktop computer with an AMD Ryzen 5 2600X CPU and 16GB RAM. In the following sections, we first evaluate the capability of iKnap, compared to both OCBC and a no-communication control condition, and then we perform ablation studies comparing iKnap to OCBC.
(a) Agent/subject count $n+m$
(b) Noise parameter $\alpha$
(c) Wall count $w$
(d) Optimizer runtime (s)
Figure 3: Capability studies for iKnap (blue), a no-communication control condition (orange), and OCBC (green). We measure makespan (s) vs. agent/subject count $n+m$, noise parameter $\alpha$, and wall count $w$. We also measure optimizer runtime (s) vs. agent/subject count $n+m$. Each configuration’s measurements are averaged across $100$ randomized trials.
(a) Communication frequency (hz)
(b) Bandwidth range
(c) Max. bandwidth proportion
(d) Prediction horizon (s)
Figure 4: Ablation studies for iKnap (blue) and OCBC (green). We measure makespan (s) vs. communication frequency (hz), range of randomly generated pairwise bandwidth values, bandwidth limit ratio $B/n$, and agent trajectory prediction horizon (s). Each configuration’s measurements are averaged across $100$ randomized trials.
IV-F Capability Studies
To assess the capability of iKnap, we measure the navigation performance with respect to the size of the communication problem, environment complexity, and observation noise. Thus, we measure makespan with respect to the combined number of agents and subjects in equal number, the number of walls in the environment, and the amount of agent sensor noise, respectively. We also measure the runtime of the communication scheme with respect to the agent and subject count. We compare iKnap’s performance to that of OCBC, as well as a no-communication baseline to exemplify the effectiveness of observation sharing between agents. By default, we configure $n=5$, $m=5$, $w=10$, and $\alpha=0.01$. For all other default values, see Section IV-G. For each parameter configuration, we simulate $100$ trials with randomized initial conditions and compute the average value.
Figure 3 displays the results of the capability experiments. These results imply that iKnap’s navigation performance scales better than OCBC and the no-communication control group with respect to scenario complexity (Fig. 3a, Fig. 3c) and under significant observational constraints (Fig. 3b). Moreover, Figure 3d displays the optimizer runtimes for iKnap and OCBC with respect to scalability. The two algorithms exhibit largely similar runtime, and even in scenarios involving $20$ robots and $20$ dynamic obstacles as subjects, iKnap poses only a $10\%$ increase in runtime over OCBC, while providing a consistent increase of $20\%-35\%$ in navigation performance across the other measured variables.
IV-G Ablation Studies
To perform ablation studies for iKnap, we measure the navigation performance with respect to communication frequency, bandwidth allowance, bandwidth variability, and utility heuristic performance. Thus, we measure makespan with respect to the frequency at which the optimal communication problem is run and its results executed, the ratio of the bandwidth limit to the number of agents, the range of randomly generated bandwidth costs associated with communications between each pair of agents, and the trajectory prediction horizon for our chosen utility heuristic. We compare iKnap’s performance to that of OCBC. By default, we configure a communication frequency of $1$ hz, a bandwidth limit ratio of $5$ units per agent, uniformly random pairwise bandwidths from $1$ to $10$ units (yielding a range of $9$ units), and a prediction horizon of $5$ seconds. For each parameter configuration, we simulate $100$ trials with randomized initial conditions and compute the average value.
Figure 4 displays the results of the ablation experiments. These results imply that iKnap utilizes allocated communication resources (Fig. 4a-c) and observational resources (Fig. 4d) more effectively than OCBC. Moreover, iKnap excels in heavily constrained scenarios.
V Conclusion
In this paper, we introduced Intelligent Knapsack (iKnap), an optimal communication scheme designed for multi-robot observation sharing in dynamic environments with bandwidth constraints. By modeling communication as graph-based multi-agent inference and framing the optimization problem as a 0/1 knapsack problem, iKnap efficiently balances the tradeoff between information value and communication cost. Our empirical analysis shows that iKnap significantly enhances navigation performance, particularly in complex scenarios, while maintaining comparable runtime to existing broadcast-based schemes. It also optimizes bandwidth and observational resource usage more effectively, especially in low-resource and high-uncertainty settings. These results suggest that iKnap facilitates robust collaboration among multi-robot teams in real-world navigation challenges. Moreover, when the subjects are human subjects, the proposed method has the potential to improve safety and efficiency in multi-robot multi-human interaction. Future research will focus on expanding iKnap’s compatibility with diverse agent behavior models and integrating learning-based approaches to construct generic communication utility functions for various scenarios. We also intend to perform real-world tests in factory environments.